OpenStudy (anonymous):

The probability that Jane's dog will catch a Frisbee when she throws it is 8 out of 12. Yesterday, she threw it 204 times. How many times did the dog catch it?

OpenStudy (anonymous):

8/12 = x/204 cross multiply 8*204= 1,632 then divide by 12, x=136 times/ 204

OpenStudy (anonymous):

Thank you.